write the following without using logarithims.
ln a - ln b + xy = 0
Thanks!

Answer by nerdybill(7384)

ln a - ln b + xy = 0
ln a - ln b = -xy
ln (a/b) = -xy
a/b = e^(-xy)
a/b = 1/e^(xy)